OpenStudy (anonymous):

help please anyone good in calculus and please explain step by step it would be really great.

OpenStudy (anonymous):

\[If F(x)=\int\limits_{0}^{x}\frac{ 1 }{ t^2+1 }dt --> then F'(x)=\]

OpenStudy (anonymous):

the derivative of the integral is the integrand

OpenStudy (anonymous):

\[F(x)=\int_a^xf(t)dt\to F'(x)=f(x)\]

OpenStudy (anonymous):

there is really no "step" here where you see a \(t\) in the expression \(\frac{1}{t^2+1}\) replace it by an \(x\)

OpenStudy (anonymous):

so its \[-\frac{ x^2 }{ x^2+1 }\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

replace \(t\) by \(x\) in \(\frac{1}{t^2+1}\) there is no minus sign there is no \(x\) in the numerator

OpenStudy (anonymous):

Fundamental Theorem of Calculus

OpenStudy (anonymous):

but the derivative is not the same

OpenStudy (anonymous):

not the same as what?

OpenStudy (anonymous):

as the original

OpenStudy (nincompoop):

they are the same

OpenStudy (anonymous):

\[F(x)=\int _0^x\sin(t)dt\] \[F'(x)=\sin(x)\]

OpenStudy (anonymous):

\[F(x)=\int_0^x e^{-\frac{t^2}{2}}dt\] \[F'(x)=e^{-\frac{x^2}{2}}\]

OpenStudy (anonymous):

ok i get now thank u all but can u check my answers for other problems because i think some of them are wrong